OpenStudy (anonymous):

You want to fence a rectangular area of your yard for a garden. Youplan to use no more than 50 ft of fencing. a. Write and graph a linear inequality that describes this situation. b. Open-Ended What are two possible sizes for a square garden?

OpenStudy (anonymous):



OpenStudy (anonymous):

A)2w + 2< = 50; is that right?

OpenStudy (anonymous):

Let x be the length & y be the width of the garden,then 2(x+y)<=50 or x+y<=25

OpenStudy (anonymous):

ill put the first one

OpenStudy (anonymous):